Undiluted solution:

% Transmission of Cobalt Chloride: 31.62%
OD of undiluted solution (OD = 2 - log10 (%T)):


Diluted solution (1:10):

% Transmission of Cobalt Chloride: 89.13%
OD of undiluted solution (OD = 2 - log10 (%T)):


Diluted solution (1:100):

% Transmission of Cobalt Chloride: 98.86%
OD of undiluted solution (OD = 2 - log10 (%T)):
Answer #1

1) OD = 2 - log10(%T) = 2 - log10(31.62) = 2 - 1.5 = 0.5

2) OD = 2 - log10(%T) = 2 - log10(89.13) = 2 - 1.95 = 0.05

3) OD = 2 - log10(%T) = 2 - log10(98.86) = 2 - 1.995 = 0.005